OpenStudy (dtan5457):

How do I factor s^2+st-2t^2? A lot of the time I just guess and check, but what's the logical way to factor this?

OpenStudy (danjs):

(s-t)(s+2t)

OpenStudy (danjs):

seems to work

OpenStudy (dtan5457):

it does. but how did you figure that out?

OpenStudy (dtan5457):

a regular trinomial you look to see what adds to the value of b and multiplies to c, what about here?

OpenStudy (danjs):

(-1)(2t + s)(t - s)

OpenStudy (dtan5457):

(2t^2 - s*t - t^2) -s^2?

OpenStudy (danjs):

(-1)(2t^2 - s*t - s^2) want -s^2 for the multiply part and -1s for the addition part \[(-1)(2t \pm~~~)(t \pm~~~~)\]

OpenStudy (dtan5457):

why -1s?

OpenStudy (danjs):

that is the coefficient on the t term

OpenStudy (dtan5457):

so leave out the t?

OpenStudy (danjs):

i put the t's in the parenthesis already, you just have to fill in the plus or minus and the s's

OpenStudy (danjs):

(-1) (2t + s)(t - s) u can see that the factors will be just 1s and 1s, then just test out where the + and - should go

OpenStudy (danjs):

+1s - 2s = -s

OpenStudy (dtan5457):

alright. i kind of get it now.

OpenStudy (dtan5457):

i'll try this the next time i get something like this

OpenStudy (dtan5457):

guess and check almost seems easier though, lol

OpenStudy (danjs):

it is just a game with the factors, you can always use the quadratic formula and figure it out if needed

OpenStudy (danjs):

yeah guess and check is pretty much it, after you list the factors

OpenStudy (dtan5457):

got it. thanks

OpenStudy (danjs):

If you use the quadratic formula for A=2 B= -s C = -s^2 you will get those factors

OpenStudy (dtan5457):

let me try

OpenStudy (danjs):

-(2t^2 - s*t - s^2) = -(2t+s)(t-s) t-s = 0 ... t = s 2t + s = 0 ....t = -1/2 s quadratic formula will give the same

OpenStudy (dtan5457):

o.

OpenStudy (danjs):

\[\frac{ s \pm \sqrt{(-s)^2 - 4 * 2 * -s^2}}{ 4 } = \frac{ s \pm \sqrt{9s^2} }{ 4 } = \frac{ s \pm 3s }{ 4 }\]

OpenStudy (danjs):

4s/4 = s -2s/4 = -1/2 s

OpenStudy (danjs):

which are the same values if you found the zeros to -(2t+s)(t-s)

OpenStudy (dtan5457):

lol wow. didn't think of it like that.